The sum of two numbers is 17. Their difference is 7. Find the numbers.
Answer by jim_thompson5910(35256)

x = first number
y = second number

The sum of two numbers is 17, so
x+y = 17

Their difference is 7, so
x-y = 7

We have this sytem of equations
x+y = 17
x-y = 7

Add up the equations to get
x+y = 17
x-y = 7
--------
2x+0y = 24

So 2x = 24 which means x = 12 after you divide both sides by 2 to isolate x.

Use x = 12 to find the other number.

x+y = 17
12+y = 17
12+y-12 = 17-12
y = 5

So the second number is y = 5

-------------------------------------------

Final Answer: the two numbers are 12 and 5

Sum: 12+5 = 17
Difference: 12-5 = 7
Answer checks out